# Taming Throughput-Latency Tradeoff in LLM Inference with Sarathi-Serve

### v1

@@ 5.2 Ablation Study
Figure 11: Overhead of chunked-prefills for Yi-34B (TP-2) shown for various operations and prompt sizes using batch size 1. To calculate the overhead, we divide prefill time with chunking by prefill time without chunking.
@@
Figure 11 shows how much overhead chunking adds in Yi-34B – on individual operators as well as the overall runtime. First, smaller chunks introduce higher overhead as shown by the gradually decreasing bar heights in Figure 11. This is expected and remains true in all cases, irrespective of the operator type or the prompt size. Second, attention and allreduce operators experience the highest overhead. The attention overhead is expected due to repeated global memory accesses. The overhead of allreduce is higher due to the fixed cost of NCCL communication collectives that get invoked more frequently with smaller chunks. Third, the linear operators experience the lowest overhead in all cases where the overhead is close to zero with chunk size 2048. Note that because the cost of linear operators dominate in transformers [71], the overall runtime overhead of chunked-prefills is close to that of the linear operators.
@@
hybrid-batching-only 0.56 1.08 4.18 1.76
chunked-prefills-only 1.04 0.34 4.86 0.38
Sarathi-Serve (combined) 0.85 0.29 3.03 0.35
Table 4: Performance impact of hybrid-batching and chunked-prefills in isolation as well as when they are used in tandem, evaluated over 128 requests for Yi-34B running on two A100s with a token budget of 1024.
@@ 7 Conclusion
We also introduce Sarathi-Serve and show that the techniques we originally proposed in Sarathi [27] (to optimize throughput alone) are equally powerful in co-optimizing throughput and latency. Sarathi-Serve chunks input prompts into smaller units of work to create stall-free schedules. This way, Sarathi-Serve can add new requests in a running batch without pausing ongoing decodes. Our evaluation shows that Sarathi-Serve improves the serving capacity of Mistral-7B by up to 2.6$\times$ on a single A100 GPU and up to 6.9$\times$ for Falcon-180B on 8 A100 GPUs.

### v3

@@ 5.4 Ablation Study
5.4.1 Overhead of chunked-prefills
Figure 14 shows how much overhead chunking adds in Yi-34B – on overall prefill runtime. As expected, smaller chunks introduce higher overhead as shown by the gradually decreasing bar heights in Figure 14. However, even with the smallest chunk size of 512, we observe a moderate overhead of at most $\sim$25%. Whereas with the larger token budget of 2048, chunked prefills have almost negligible overhead.
Figure 14: Overhead of chunked-prefills in prefill computation for Yi-34B (TP-2) normalized to the cost of no-chunking, shown for various prompt lengths using chunk lengths of 512, 1024 and 2048.
@@ 5.4.1 Overhead of chunked-prefills
hybrid-batching-only 0.53 0.68 3.78 1.38
chunked-prefills-only 1.04 0.17 5.38 0.20
Sarathi-Serve (combined) 0.76 0.14 3.90 0.17
Table 4: TTFT and TBT latency measured in seconds for hybrid-batching and chunked-prefills used in isolation as well as when they are used in tandem, evaluated over 128 requests for Yi-34B running on two A100s with a token budget of 1024. By using both hybrid-batching and chunked-prefills, Sarathi-Serve is able to lower both TTFT and TBT.
5.4.2 Impact of individual techniques
@@ 7 Conclusion
[…]
8 Acknowledgement
We would like to thank OSDI reviewers and our shepherd for their insightful feedback. This research is partly supported by GT Cloud Hub, under the auspices of the Institute for Data Engineering and Science (IDEaS), with funding from Microsoft, and the Center for Research into Novel Compute Hierarchies (CRNCH) at Georgia Tech.
Appendix A Artifact Appendix
Abstract
Our open source artifact is available on GitHub. This repository contains our implementation of Sarathi-Serve as well as the harnesses and scripts for running and plotting the experiments described in this paper.
This repository originally started as a fork of the vLLM project. Sarathi-Serve is a lightweight high-performance research prototype and doesn’t have complete feature parity with open-source vLLM. We have only retained the most critical features and adopted the codebase for faster research iterations.
Scope
This artifact allows the readers to validate the claims made in the Sarathi-Serve paper (the figures) and provides a means to replicate the experiments described. The artifact can be used to set up the necessary environment, execute the main results, and perform microbenchmarks, thus providing a comprehensive understanding of the key claims in Sarathi-Serve.
Contents
The repository is structured as follows, the primary source code for the system is contained in directory /sarathi. The implementations for custom CUDA kernels are within the /csrc directory. All the scripts to reproduce the experiments are in /osdi-experiments and finally, the trace files used for the experiments are stored in /data.
Hosting
You can obtain our artifacts from GitHub: GitHub. The main branch of the Github repository is actively updated, but we will maintain clear and accessible instructions about our artifacts in an easily identifiable README file. All the detailed instructions and README files to reproduce the experiments in the OSDI paper are available in the branch osdi-sarathi-serve.
Requirements
Sarathi-Serve has been tested with CUDA 12.1 on A100 and A40 GPUs. The specific GPU SKUs on which the experiments were performed and the parallelism strategies used are clearly explained in the README corresponding to the figures in the artifact, for ease of reproducibility.